# CiFlow: Dataflow Analysis and Optimization of Key Switching for Homomorphic Encryption

### v2

@@ Abstract
In this paper, we present a novel approach to improve key-switching performance by rigorously analyzing its dataflow. Our primary goal is to optimize data reuse with limited on-chip memory to minimize off-chip data movement. We introduce three distinct dataflows: Max-Parallel (MP), Digit-Centric (DC), and Output-Centric (OC), each with unique scheduling approaches for key-switching computations. Through our analysis, we show how our proposed Output-Centric technique can effectively reuse data by significantly lowering the intermediate key-switching working set and alleviating the need for massive off-chip bandwidth. We thoroughly evaluate the three dataflows using the RPU, a recently published vector processor tailored for ring processing algorithms, which includes HE. This evaluation considers sweeps of bandwidth and computational throughput, and whether keys are buffered on-chip or streamed. With OC, we demonstrate up to 4.16$\times$ speedup over the MP dataflow and show how OC can save 16$\times$ on-chip SRAM by streaming keys for minimal performance penalty.
@@ I Introduction
[…]
The slowdown has been addressed by a growing body of prior work on hardware acceleration for HE. Many solutions now exist that consider fixed-function pipelines [9], vector architectures [10, 13, 14], and tiled architectures [15, 16]. The architectural approaches taken differ, but the building blocks are the same: very large on-chip memories and high off-chip bandwidth. For example, CraterLake [13]) uses 256 MB of on-chip SRAM and assumes two HBM2e PHYs for a total bandwidth of 1TB/s. BTS [15] and ARK [16] both have 512 MB of on-chip SRAM and assume 1TB/s of memory bandwidth. The large on-chip SRAMs and multiple PHYs result in large chips, adding the cost of the advanced memory technology results in expensive solutions.
Our hypothesis is that by optimizing HE dataflow, we can capture on-chip reuse with far less SRAM and simultaneously reduce the off-chip bandwidth requirements. To understand the potential of optimizing dataflow, we will deep dive into the hybrid key-switching algorithm (HKS) [17]. HKS is the core computation of HE and highly complex. The use of HKS in HE is widespread; it is called after each rotation and homomorphic multiplication and is heavily used in bootstrapping. For example, recent work has reported that a single HE ResNet-20 inference takes 3,306 rotations [18], and prior research has further shown that HKS dominates HE’s runtime [19, 16, 20], and can be up to 70% for private neural inference [18].
@@
To understand and optimize HKS, we propose three distinct dataflows: Max-Parallel (MP), Digit-Centric (DC), and Output-Centric (OC). Our key insight is that with the OC dataflow, the intermediate state of HKS can be significantly compressed while still maintaining high parallelism to utilize computational units. To demonstrate the potential of OC, we implement five parameterizations of HKS taken from recently published work following each of the three presented dataflows on a recent vector HE accelerator [14]. Using previously validated simulation infrastructure, we rigorously evaluate the dataflows. First, we assume a large on-chip SRAM (i.e., 512 MB) as in prior work and assume it can be used to buffer all $\mathbf{evks}$ on-chip while leaving a smaller fraction (i.e., 32MB) for data (i.e., inputs and intermediates). We find that OC consistently outperforms other dataflows and can deliver up to 4.16$\times$ speedup over MP using the same bandwidth. Next, given that $\mathbf{evks}$ have no reuse within HKS, we elect to stream them on-chip and reserve a fraction of off-chip bandwidth for them. Here again OC performs well; by streaming $\mathbf{evks}$ we can reduce the on-chip SRAM requirements by 16$\times$ while still having 3.3$\times$ less bandwidth compared to MP using 64GB/s bandwidth and storing $\mathbf{evks}$ on-chip.
In summary our contributions are:
@@
    We propose three different dataflows for HKS algorithm, named Max-Parallel, Digit-Centric, and Output-Centric. Through our analysis we demonstrate how strategically scheduling instructions and efficiently reusing loaded and generated data by OC dataflow can lead to substantial off-chip bandwidth saving.
@@
    We evaluate two scenarios for handling $\mathbf{evks}$: streaming from off-chip and on-chip storage with a larger memory. Streaming $\mathbf{evks}$ reduces SRAM area by 16$\times$ and OC saves 3.3$\times$ bandwidth over the MP baseline.
@@
    We evaluate the performance of dataflows across varing off-chip bandwidths and computational throughput to understand the bandwidth-compute trade-off. We find OC to be highly effective, matching naive MP HKS implementation with significant bandwidth saving. Furthermore, increasing accelerator throughput enhances performance even further.
@@ II Background
In this section, we briefly introduce the CKKS HE scheme [5], including its key parameters and operations, using terminology from ARK[16] and Castro [19] when applicable. Unlike prior HE schemes, BGV [2] and BFV[3, 4], CKKS operates on vectors of real or complex numbers, rather than integers. To perform these operations, a vector message, $\mathbf{m}$, is first encoded into a plaintext polynomial $R_{Q}=\mathbb{Z}_{Q}[X]/(X^{N}+1)$, represented as $[\mathtt{P}]$. In essence, this is another vector of length $N$, with each element being an integer no larger than Q. Depending on a problem’s complexity, the value of $N$ typically ranges from $2^{11}$ to $2^{17}$, while $Q$ can range from hundreds to thousands of bits. For a polynomial with $N$ coefficients, a vector message must be of length $n\leq\nicefrac{{N}}{{2}}$.
@@
CKKS supports several arithmetic operations like addition and multiplications between plaintext polynomials or ciphertexts. Additions and multiplications are simple point-wise operations between coefficients and therefore require the same number of levels in both operands. CKKS also supports rotations that cyclically rotates elements within the vector message by a specified amount, $r$. Notably, indexing vector elements is not efficiently supported and therefore ciphertext rotations are the primary way of computing fully connected layers and convolutions in neural networks.
@@
III Hybrid Key Switch
In this section, we will describe the Hybrid key-switching (HKS) algorithm, a crucial step after performing ciphertext rotations and ciphertext-ciphertext multiplications. Both operations convert a ciphertext that is encrypted by a secret key, $s$, into a new ciphertext only decryptable by a new secret key, $s^{\prime}$. In general, to continue computation, the ciphertext must be returned to a form decryptable by the original key, $s$. The act of “switching” the keys from $s’$ ’ back to $s$ is known as key-switching. This is achieved by multiplying the ciphertext $[\![\mathbf{m}]\!]_{s^{\prime}}$, with a special evaluation key, $\mathbf{evk}$, which re-encrypts it from $s^{\prime}$ to $s$. However, this process comes with an exceedingly high level of noise, which keeps growing with consecutive use of the key-switching algorithm. Techniques proposed by Fan [4] and later by Han [17] address this issue by performing the $\mathbf{evk}$ multiplication at a higher modulus. Specifically, from Table I, this multiplication is performed on polynomials, $[\mathtt{P}]_{\mathcal{D}}$ (modulo $PQ_{\ell}$), for some large integer $P$ rather than just $Q_{\ell}$. While this substantially reduces noise growth, it increases HE computational complexity by adding complex NTTs and RNS basis extensions.
HKS [17] is a generalization of existing key-switching techniques that let the user trade off the complexity of key-switching with the size of the $\mathbf{evk}$. It has been widely adopted both in software implementations [23, 24] and in recent HE accelerators [13, 15, 16, 20] yet the aforementioned trade off, especially in hardware, is still not well understood.
@@ III-A Overview
Broadly, HKS is composed of two phases that we denote as ModUp and ModDown. The ModUp phase can be broken down into five consecutive stages, P1-P5 and the ModDown phase into four, P1-P4. We use terms such as ModDown_Pi to refer to the i’th stage of the ModDown phase. These stages can be seen in Figure 2. We loosely adopt this terminology from Han [17] to intuitively represent that the first phase of HKS is geared towards extending the RNS base from $Q_{\ell}$ to $PQ_{\ell}$, whereas the second phase is focused on reducing the modulus back from $PQ_{\ell}$ to $Q_{\ell}$.
We consider an input polynomial as a matrix of size $(N\times\ell)$ which changes shape throughout the key-switching process, depending on the chosen parameters. Figure 2 represents the HKS dataflow for a specific parameter set, however, the core stages of HKS remain the same across different parameter choices. In Section V-B, we explore a wider range of parameter sets and note that the insights we gain from analyzing this specific case remain applicable. We highlight a single tower of the input polynomial in red in Figure 2, with the widths of subsequent stages accurately reflecting the change in sizes of the original $(N\times\ell)$ input polynomial. For example, a preliminary step in HKS decomposes the input polynomial into $dnum$ digits, each of size roughly, $\alpha=\lceil\nicefrac{{(L+1)}}{{dnum}}\rceil$, reshaping the input matrix to $(N\times\ell)$. In Figure 2, for which $dnum=3$, the input polynomial is decomposed into three different colored digits, with each digit being $11$ towers wide.
@@ III-B ModUp
[…]
@@ III-C ModDown
[…]
IV HKS Dataflows
We propose three dataflows for the HKS algorithm: Max-Parallel (MP), Digit-Centric (DC), and Output-Centric (OC). These dataflows differ in their sequence of instructions, reusing the loaded and computed data, intermediate data generation, and off-chip memory interaction. Assuming unlimited on-chip memory, the performance of these dataflows would be identical, as all inputs and intermediate data can be stored on-chip. Our goal is to show by changing the sequence of operations we can have the same performance using a smaller on-chip memory and less off-chip bandwidth. In this section, we provide a detailed explanation of the principles and strategies behind each dataflow. Later we evaluate their performance by varying bandwidth and computational throughput. Our analysis highlights how the dataflow will help to save memory and bandwidth resources.
@@ IV-A Max-Parallel Dataflow (MP)
This dataflow is designed to prioritize kernel parallelism at all costs. Each tower is loaded onto the on-chip memory for a single operation, and each operation is executed on all input towers sequentially before processing the next step. However, this approach comes with a drawback. Specifically, in ModUp_P2 and ModDown_P2, the generated intermediate data after BConv becomes extremely large, even though only one of the BConv output towers is required for calculating each output tower. For BTS3, at least 675MB on-chip memory is required to prevent excessive load and stores to the off-chip.
@@ IV-B Digit-Centric Dataflow (DC)
This dataflow adopts a “one-digit-at-a-time” approach, where each digit is loaded onto the chip, and all possible calculations involving that digit are performed before moving on to the next digit. As illustrated in Figure 2(b), all P1 to P5 calculations for a single digit are done in sequence, maximizing data reuse. However, BConv still expands the digit from $\alpha$ to $\beta$ towers, generating a partial product of the output that can either be stored on-chip for later reduction to save bandwidth or sent off-chip to minimize on-chip memory requirement. Once ModUp is completed, ModDown follows the same approach. For BTS3, which is the largest benchmark, DC requires 255MB of on-chip memory, which is 62% less compared to MP dataflow.
@@ IV-C Output-Centric Dataflow (OC)
Finally, we propose Output-Centric (OC) dataflow, which is optimized for data movement and on-chip storage, leveraging insights from the first two solutions. OC focuses on computing one output tower at a time, efficiently utilizing on-chip resources. In this dataflow, the ModUp stage has two sections: Section1 computes output towers in modulo $Q$, with $dnum-1$ digits going through ModUp_P2 & ModUp_P3 step and one tower from the last digit bypassed through ModUp_P2 & ModUp_P3. Section2 computes output towers in mod $P$, requiring all digits to pass through ModUp_P2 & ModUp_P3 for a single output calculation.
In both sections, the computation is optimized to minimize the on-chip memory requirement. Since only one output tower is calculated at a time, the entire computation of ModUp_P2 & ModUp_P5 is unnecessary. Instead, a single ModUp_P2 output tower is calculated and multiplied with the $\mathbf{evk}$, reducing memory requirements and off-chip data communication. Moreover, for ModUp_P5, only one partial tower is calculated at a time, allowing them to be aggregated before being stored back. In Section2, where all digits are required for calculating a single output tower, we have used the following strategy to manage the on-chip memory limitation. Since the INTT of the first $dnum-1$ digits are already on-chip, so we compute the partial sum with those digits and then the final digit is loaded to compute the last partial sum and the final output towers. This approach reduces off-chip memory interaction and on-chip memory requirement.
ModDown has the same approach as ModUp, loading all towers related to $[\mathtt{P}]_{\mathcal{C}}$ on-chip. Calculating one output tower at a time, eliminates the expansion of ModUp_P2, enhancing the efficiency of HKS with a small on-chip memory.

### v4

@@ Abstract
In this paper, we present a novel approach to improve key-switching performance by rigorously analyzing its dataflow. Our primary goal is to optimize data reuse with limited on-chip memory to minimize off-chip data movement. We introduce three distinct dataflows: Max-Parallel (MP), Digit-Centric (DC), and Output-Centric (OC), each with unique scheduling approaches for key-switching computations. Through our analysis, we show how our proposed Output-Centric technique can effectively reuse data by significantly lowering the intermediate key-switching working set and alleviating the need for massive off-chip bandwidth. We thoroughly evaluate the three dataflows using the RPU, a recently published vector processor tailored for ring processing algorithms, which includes HE. This evaluation considers sweeps of bandwidth and computational throughput, and whether keys are buffered on-chip or streamed. With OC, we demonstrate up to 4.16$\times$ speedup over the MP dataflow and show how OC can save 12.25$\times$ on-chip SRAM by streaming keys for minimal performance penalty.
@@ I Introduction
Many HE schemes (e.g., BGV [2], BFV [3, 4], and CKKS [5]) and software implementations (e.g., OpenFHE [6], Lattigo [7], SEAL [8]) now exist and have made steady performance improvements over the years. However, practical deployment of HE is still limited by significant performance overheads, attributed to large data sizes, complex operations, and additional functions to process. When encrypting and processing HE ciphertexts, the data becomes significantly larger than plaintext, forming large ciphertext vectors of up to 217 elements. Here, each plaintext element is represented by hundreds to thousands of bits, and special precomputed keys ($\mathbf{evks}$) are needed for certain operations (i.e., multiplication and rotation) that can be upwards of hundreds of MBs. This puts substantial pressure on the memory system. Next, HE requires complex modular arithmetic for processing each element, which is not natively supported by mainstream commercial processors. Finally, HE operations are not one-to-one conversions of plaintext operators; they involve multiple additional functions for performance (e.g., Number Theoretic Transforms (NTT)) and correctness before and after the actual operation (e.g., multiplication and rotation). These factors collectively overwhelm commodity hardware, and prior work has consistently reported 4-6 orders of magnitude slowdown compared to plaintext [9, 10, 11, 12].
The slowdown has been addressed by a growing body of work on hardware acceleration for HE. Many solutions now exist that consider fixed-function pipelines [9, 13], vector architectures [10, 14, 15], and tiled architectures [16, 17]. The architectural approaches taken differ, but the building blocks are the same: very large on-chip memories and high off-chip bandwidth. For example, CraterLake [14]) uses 256 MB of on-chip SRAM and assumes two HBM2e PHYs for a total bandwidth of 1TB/s. BTS [16] and ARK [17] both have 512 MB of on-chip SRAM and assume 1TB/s of memory bandwidth. The large on-chip SRAMs and multiple PHYs result in large chips, adding the cost of the advanced memory technology results in expensive solutions.
Our hypothesis is that by optimizing HE dataflow, we can capture on-chip reuse with far less SRAM and simultaneously reduce the off-chip bandwidth requirements. To understand the potential of optimizing dataflow, we deeply analyze the hybrid key-switching algorithm (HKS) [18]. HKS is the core computation of HE and highly complex. The use of HKS in HE is widespread; it is called after each rotation and homomorphic multiplication and is heavily used in bootstrapping. For example, recent work has reported that a single HE ResNet-20 inference takes 3,306 rotations [19], and prior research has further shown that HKS dominates HE’s runtime [20, 17, 21], and can be up to 70% for private neural inference [19].
@@
To understand and optimize HKS, we propose three distinct dataflows: Max-Parallel (MP), Digit-Centric (DC), and Output-Centric (OC). Our key insight is that with the OC dataflow, the intermediate state of HKS can be significantly compressed while maintaining high parallelism to utilize computational units. To demonstrate OC’s potential, we implement five parameterizations of HKS taken from recent work following each of the three presented dataflows on a recent vector HE accelerator [15]. Using previously validated simulation infrastructure, we rigorously evaluate the dataflows. First, we assume a large on-chip SRAM (i.e., 392 MB), which is sufficient to buffer all $\mathbf{evks}$ on-chip while reserving 32MB for data (i.e., inputs and intermediates). We find that OC consistently outperforms other dataflows and can deliver up to 4.16$\times$ speedup over MP using the same bandwidth. Next, given that $\mathbf{evks}$ have no reuse within HKS, we elect to stream them on-chip and reserve a fraction of off-chip bandwidth for them. Here again, OC performs well. With our OC dataflow, we can stream $\mathbf{evks}$ to reduce on-chip SRAM by 12.25$\times$ and bandwidth by 3.3$\times$, while achieving the same performance as an MP on-chip implementation.
In summary, our contributions are:
@@
    We propose three different dataflows for HKS algorithm, named Max-Parallel, Digit-Centric, and Output-Centric. Through our analysis we demonstrate how strategically scheduling instructions and efficiently reusing loaded and generated data by the OC dataflow can lead to substantial off-chip bandwidth saving.
@@
    We evaluate two scenarios for handling $\mathbf{evks}$: streaming from off-chip and on-chip storage with a larger memory. Streaming $\mathbf{evks}$ reduces SRAM area by 12.25$\times$ and OC saves 3.3$\times$ bandwidth over the MP baseline.
@@
    We evaluate the performance of dataflows across different off-chip bandwidths and computational throughput to understand the bandwidth-compute trade-off. We find OC to be highly effective, matching naive MP HKS implementation with significant bandwidth saving. Furthermore, increasing accelerator throughput enhances performance even further.
@@ II Background
In this section we briefly introduce the CKKS HE scheme [5], including its key parameters and operations, using terminology from ARK[17] and Castro [20] when applicable. Unlike prior HE schemes, BGV [2] and BFV[3, 4], CKKS operates on vectors of real or complex numbers, rather than integers. To perform these operations, a vector message, $\mathbf{m}$, is first encoded into a plaintext polynomial $R_{Q}=\mathbb{Z}_{Q}[X]/(X^{N}+1)$, represented as $[\mathtt{P}]$. In essence, this is another vector of length $N$, with each element being an integer no larger than Q. Depending on a problem’s complexity, the value of $N$ typically ranges from $2^{11}$ to $2^{17}$, while $Q$ can range from hundreds to thousands of bits. For a polynomial with $N$ coefficients, a vector message must be of length $n\leq\nicefrac{{N}}{{2}}$.
@@
CKKS supports several arithmetic operations like addition and multiplications between plaintext polynomials or ciphertexts. Additions and multiplications are simple point-wise operations between coefficients and therefore require the same number of levels in both operands. CKKS also supports rotations that cyclically rotate elements within the vector message by a specified amount, $r$. Notably, indexing vector elements is not efficiently supported and therefore ciphertext rotations are the primary way of computing fully connected layers and convolutions in neural networks.
@@
III Hybrid Key-Switching
In this section, we describe the Hybrid key-switching (HKS) algorithm, a crucial step after performing ciphertext rotations and ciphertext-ciphertext multiplications. Both operations convert a ciphertext encrypted by a secret key, $s$, into a new ciphertext only decryptable by a new secret key, $s^{\prime}$. To continue computation, the ciphertext must be returned to a form decryptable by the original key, $s$. This process, known as key-switching, involves multiplying the ciphertext $[\![\mathbf{m}]\!]_{s^{\prime}}$ with a special evaluation key, $\mathbf{evk}$, which re-encrypts it from $s^{\prime}$ to $s$. However, this process introduces significant noise growth, which can be managed using techniques proposed by Fan [4] and later by Han [18]. These techniques involve performing the $\mathbf{evk}$ multiplication at a higher modulus, reducing noise growth but increasing HE computational complexity with complex NTTs and RNS basis extensions.
HKS [18] is a generalization of existing key-switching techniques that let the user trade off the complexity of key-switching with the size of the $\mathbf{evk}$. It has been widely adopted both in software implementations [24, 25] and in recent HE accelerators [14, 16, 17, 21] yet the aforementioned trade-off, especially in hardware, is still not well understood.
@@ III-A Overview
Broadly, HKS is composed of two phases that we denote as ModUp and ModDown. The ModUp phase can be broken down into five consecutive stages, P1-P5, and the ModDown phase into four, P1-P4. We use terms such as ModUp/Down_Pi to refer to the i’th stage of the ModUp/Down phase. These stages can be seen in Figure 2. We loosely adopt this terminology from Han [18] to intuitively represent that the first phase of HKS is geared towards extending the RNS base from $Q_{\ell}$ to $PQ_{\ell}$, whereas the second phase is focused on reducing the modulus back from $PQ_{\ell}$ to $Q_{\ell}$.
We consider an input polynomial as a matrix of size $(N\times\ell)$, which changes shape throughout the key-switching process, depending on the chosen parameters. Figure 2 represents the HKS dataflow for a specific parameter set; the core stages of HKS remain the same across different parameter choices. In Section V-B, we explore a wider range of parameter sets and note that the insights we gain from analyzing this specific case remain applicable. We highlight a single tower of the input polynomial in red in Figure 2, with the widths of subsequent stages accurately reflecting the change in sizes of the original $(N\times\ell)$ input polynomial. For example, a preliminary step in HKS decomposes the input polynomial into $dnum$ digits, each of size roughly, $\alpha=\lceil\nicefrac{{(L+1)}}{{dnum}}\rceil$, reshaping the input matrix to $(N\times\ell)$. In Figure 2, for which $dnum=3$, the input polynomial is decomposed into three different colored digits, with each digit being $11$ towers wide.
@@ III-B ModUp
$\mathbf{(P1)}\hskip 1.99997pt\textit{INTT}$: Following the initial digit decomposition step, each tower must undergo an INTT, which converts the polynomial from the evaluation domain to the coefficient domain. This conversion has a complexity of $\mathcal{O}(N\log N)$, similar to the FFT, and is applied separately to all $\ell$ towers.
$\mathbf{(P2)}\hskip 1.99997pt\textit{BConv}$: Each digit, now in the coefficient domain, goes through a basis conversion to change the set of primes representing a polynomial. The number of towers is extended from $\alpha$ to $\beta$, where $\beta=\ell+K-\alpha$, and K is the number of towers in moduli $P$. The number of modular multiplications in this stage is roughly $N\times\alpha\times\beta$ for each of the $dnum$ digits.
$\mathbf{(P3)}\hskip 1.99997pt\textit{NTT}$: After basis extension, the digits once again return to the evaluation domain through the NTT. The computational complexity, similar to the INTT stage, is $\mathcal{O}(N\log N)$, performed independently for $\beta\times dnum$ towers. This output is concatenated with the original digit to form a new polynomial modulo $PQ_{\ell}$. Our new matrix is now of size $dnum\times N\times(\ell+K)$.
@@
$\mathbf{(P5)}\hskip 1.99997pt\textit{Reduce}$: The last stage in the ModUp phase sums each digit’s output from ${P4}$ into one final matrix of size $2N(\ell+K)$.
@@ III-C ModDown
We now have two polynomials, each of size $N\times(\ell+K)$ that must be reduced to their original size of $N\times\ell$ for further computation. This process begins by taking the last $K$ towers of each polynomial and performing a similar series of $\textit{INTT}\rightarrow\textit{BConv}\rightarrow NTT$ operations. Here, each BConv operation converts the number of towers from $K$ to $\ell$. The runtime complexity of these three stages is $2K\times N\log N$, $2N\times K\times\ell$, and $2\ell\times N\log N$, respectively. These polynomials go through one final scalar-tower multiplication and summation to finish out the key-switching process.
IV CiFlow: A Taxonomy and Analysis
of HKS Dataflow
We propose three dataflows for the HKS algorithm: Max-Parallel (MP), Digit-Centric (DC), and Output-Centric (OC). These dataflows differ in their sequence of instructions, reuse of loaded and computed data, intermediate data generation, and off-chip memory interaction. Assuming unlimited on-chip memory, the performance gap between these dataflows would decrease significantly. This is because all inputs and intermediate data could be stored on-chip, making MP even more advantageous than DC and OC due to its highly parallel nature. (E.g., this is the dataflow Cheetah [9] used to demonstrate the FHE overhead could be overcome with large chips.) Our goal is to show that by changing the sequence of operations we can have the same performance using a smaller on-chip memory and less off-chip bandwidth. In this section, we provide a detailed explanation of the principles and strategies behind each dataflow. Later, we evaluate their performance by varying bandwidth and computational throughput. Our analysis highlights how dataflow optimizations help to save memory and bandwidth resources.
@@ IV-A Max-Parallel Dataflow (MP)
This dataflow is designed to prioritize kernel parallelism at all costs. Each tower is loaded onto the on-chip memory for a single operation, and each operation is executed on all input towers sequentially before processing the next step. However, this approach comes with a drawback. Specifically, in ModUp_P2 and ModDown_P2, the generated intermediate data after BConv becomes extremely large, even though only one of the BConv output towers is required for calculating each output tower. For BTS3, at least 675MB of on-chip memory is required to prevent excessive load and stores to the off-chip. MP was used by prior work (e.g., Cheetah [9], and HEAX [13]), and we use it as the baseline HKS implementation.
@@ IV-B Digit-Centric Dataflow (DC)
This dataflow adopts a “one-digit-at-a-time” approach, where each digit is loaded onto the chip and all possible calculations involving that digit are performed before moving on to the next digit. As illustrated in Figure 2(b), all P1 to P5 calculations for a single digit are done in sequence, maximizing data reuse. The blue frames in Figure 2 show the intermediate data generated from processing a single digit in the ModUp section, which is used for calculating a partial product of the ModUp_P5 step. As depicted in Figure 2, BConv still expands the digit from $\alpha$ to $\beta$ towers, generating a partial product of the output that can either be stored on-chip for later reduction to save bandwidth or sent off-chip to minimize on-chip memory requirements. Once ModUp is completed, ModDown follows the same approach. For BTS3, which is the largest benchmark, DC requires 255MB of on-chip memory, which is 62% less compared to MP dataflow.
@@ IV-C Output-Centric Dataflow (OC)
Finally, we propose Output-Centric (OC) dataflow, which is optimized for data movement and on-chip storage, leveraging insights from the first two solutions. OC focuses on computing one output tower at a time, efficiently utilizing on-chip resources. In this dataflow, the ModUp stage has two sections: Section1 computes ModUp output towers in modulo $Q$, with $dnum-1$ digits going through ModUp_P2 & ModUp_P3 step and one tower from the last digit being bypassed through ModUp_P2 & ModUp_P3. As shown by the red towers in Figure 2, to generate the first output tower, the first digit is bypassed, while the other two digits must pass through the ModUp_P2 stage. Section2 computes ModUp output towers in mod $P$, requiring all digits to pass through ModUp_P2 & ModUp_P3 for a single output calculation.
In both sections, the computation is optimized to minimize on-chip memory requirements. Since only one output tower is calculated at a time, the entire computation of ModUp_P2 & ModUp_P4 is unnecessary. The red towers in Figure 2 represent the computations needed to generate one output tower, with ModUp being in Section1. As shown, there is no need to do all calculations of ModUp_P2 & ModUp_P4, minimizing the memory requirement and off-chip data communication. Additionally, for ModUp_P5, only one partial tower is calculated at a time, allowing them to be on-chip for accumulation and only store back the accumulation result.
In Section2, where all digits are required for calculating a single output tower in ModUp, we have used the following strategy to manage the on-chip memory limitation. Since the INTT of the first $dnum-1$ digits are already on-chip, we compute the partial sum with those digits and then the final digit is loaded to compute the last partial sum and the final output towers. This approach reduces off-chip memory interaction and on-chip memory requirement.
ModDown has the same approach as ModUp, loading all towers related to $[\mathtt{P}]_{\mathcal{C}}$ on-chip. Calculating one output tower at a time eliminates the expansion of ModDown_P2, enhancing the efficiency of HKS with a small on-chip memory.